Director's Time

Observers: B.Rothberg, O. Kuhn (Tucson)

ISA: B.Rothberg (Tucson)
Telescope Operator: S. Allanson (LBTO)
AO Support: J.Christou
Telescope Support: K.Newton/J.Little (LBTO)
Software Support: C.Baddick
Instrument Support: J. Power (LBTO)
Engineering Support: M.Gusick

Summary

Delayed start due to LUCI-2 re-installation and testing on scope.

Starting the night with LBC-R/B in beam - CV Recovery Program

Details

UT 01:12 - Problem bringing up LBC-R - error "filter systems encountered a problem: has problems on power line(5) [src/channel/channel.c:530]
- related to IT 5369 - trying lbckill, deleteing power.conf, lbcstart - seemed to recover OK - taking checkout biases to test.

UT 01:26 - Red HK problem (pressure). Issued LBCKILL/LBCSTART - OK now

UT 01:36 - LUCI2 work done - telescope reconfiguring for LBCs

UT 02:09 - Taking g,r SKYFLATS PA=0 (last frames 6k, 5k)

UT 02:13 - Taking g,r SKYFLATS PA=180 8-10k

UT 02:18 - Switching to PA=180 B,I skyflats

UT 02:20 - Taking B,I SKYFLATS PA=180 scaled x13 (B is a little low ~7-8k)

UT 02:25 - Taking B,I Skyflats again to improve S/N for B (scale is now 20x) - B is 5-7k again

UT 02:31 - Taking B,I Skyflats, PA=0 Scale=50x 4-6k B, 6k I

UT 02:42 - Final set B,I Skyflats PA=0 scale=100x - completed 3 each

UT 02:52 - Moving to Co-pointing field to focus - Running WRS

UT 03:09 - DIMM 0.87"

UT 03:11 - Stopped WRS - going back to DOFPIA

===================================

UT 03:22 - Converged Blue 1.33" Red 0.62"

UT 03:22 - Co-pointing check

UT 03:25 - Starting C.Veillet Program OB2new.ob

UT 03:27 - DIMM 0.82"

UT 03:33 - Completed OB2new.ob

UT 03:34 - Going to OB3new.ob field to re-run DOFPIA (blue looks elongated)

UT 03:37 - Running DOFPIA

UT 03:41 - Converged 0.85" Blue 0.51" Red

UT 03:42 - Starting OB3new.ob

UT 03:50 - Finished OB3new.ob

===================================

UT 03:51 - Moving to RU-152 std field (Uspec,U,B,g-sloan,V,I,r-sloan 1 sec exposures) - RU152_Standard_new airmass 1.309

UT 04:04 - Moving to NGC 2655 co-pointing (will attempt at least one more OB due to FANTASMAGORICAL SEEING!!!)

UT 04:07 - Running DOFPIA (mirror temp +8.2C) 0.87" blue, 0.47" red

UT 04:13 - Starting NGC_2655_GC_OB5.ob

UT 04:20 - 0.72" at U-spec

UT 04:47 - Finished OB - running DOFPIA - Blue 1.68" Red 0.68"

UT 04:52 - Running NGC_2655_GC_OB6.ob

UT 04:55 (and 4:45) - Tracker with 0s on blue - stopping OB after 9/30 (Uspec completed)

UT 05:12 - Restarting NGC_2655_GC_OB6.ob

UT 05:39 - DIMM 0.75"

UT 05:47 - Completed OB

UT 05:48 - Moving to RU-152 - Airmass 1.919-2.05

Hit the warning for focus on the Right Side

Finished with LBC for now

===================================

UT 06:00 - Switching to LUCI

UT 06:15 - Chandra AGN Program - Rothberg

First Target SDSSJ1036+0221 - Using 1" slit (unlike previous double AGN candidates) because this is a coalesced object.

UT 06:25 - Acquiring Target - GCS seeing 0.45"-0.5"

UT 06:36 - Starting spectroscopic integration -airmass 1.254

UT 07:00 - Moving to Telluric A0V - HIP 56736 - airmass 1.054

UT 07:16 - Starting Telluric spectroscopic integration

UT 07:21 - Telluric complete

UT 07:22 - Moving to SDSSJ12243+555 - Bulgless AGN candidate

UT 07:26 - Guidestar did not appear - checking the pointing

UT 07:28 - resent preset - guidestar found

UT 07:36 - Starting spectroscopic integration - Airmass 1.090

UT 07:46 - GCS seeing 0.46"

UT 07:59 - Added two more 5min exposures

UT 08:08 - Guider seeing ~0.5"

UT 08:10 - Complete

===================================

UT 08:15 - Slew to 4c+29.30 (had to check pointing). Acquisition script running...detxy offsets being sent: 04.86, -1.84
  • 30,31 are sky,obj. 32 is slit. 33 is confirmatory thru-slit image.

UT 08:35 - Start observation. Seeing is 0.9-1" even at this airmass 2.1 (taking advantage of the rigid secondary being mounted...)

UT 09:20 - Observation ended. Seeing still 0.9-1" on guider. Airmass 3

UT 09:21 - mirror, N3.75, Ks, FlexComp on. Slit image. (42-44). Object should be 48" below 1024. Field image (45-47). Field image at +48" (48-50). Forgot to change save mode from Normal

UT 09:30 - slew to telluric HIP 46625 F8V ... 4.01, -1.29 offsets applied. (53 is confirmatory thru-slit)

UT 09:40 - telluric spec started. airmass 2.37. ended at airmass 2.54

===================================

UT 09:55 - Switching back to LBC for C.Veillet recovery program - taking LBC Biases during changeover

UT 10:06 - CoPointing OB for DOFPIA focus

UT 10:10 - Running DOFPIA

UT 10:15 - Converged to 0.67" Blue 0.45" Red

UT 10:16 - Running Co-Pointing

UT 10:22 - Running OB4 #1

UT 10:38 - OB4 #1 done

UT 10:42 - Running DOFPIA

UT 10:46 - Running OB4 #2

UT 11:03 - Running OB4 #3

UT 11:20 - Completed OB4

UT 11:20 - Doing DOFPIA on OB5 field

UT 11:22 - Running DOFPIA

UT 11:26 - Converged to 0.97" Blue 0.54" Red

UT 11:27 - Starting OB5 #1 - seems ok - gradient of 30k to 20k in red, 10k-15k in blue

UT 11:57 - OB5 #1 done

UT 11:57 - DOFPIA

UT 11:58 - Blue chip trackers have 0 values again - once DOFPIA completes, will repower tracker

UT 12:06 - Starting OB5 #2

UT 12:22 - Blue 121350 Saturated

UT 12:25 - Red 121922 Saturated

UT 12:28 - Powering off LBC, switching back to LUCI to attempt one last telluric standard

UT 12:35 - Slew to telluric. 60 is image, 61 slit

UT 12:51 - shift image .. offset detxy 6.07 -0.06. Lost guiding, offset in track mode.

UT 12:54 - start observation. No spectrum seen, even when sky subtracted.

UT 12:00 - closing. Going to zenith for calibrations.

UT 01:08 - Starting LUCI Calibrations - 1" flats and arcs

UT 01:21 - 64-75 bad - counts too high - changing from 20sec to 8sec

UT 01:29 - ran arcs 2x

UT 13:40 - start 0.5" slit flats. ...turned lamp on for a test. 105-107 had lamp on, when it should have been off. 108-112 are darks, 113-117 are lamp on flats.

UT 13:58: - 0.5" comps. 118 is test Ar... last comp is 126

UT 14:04 DONE
Topic revision: r4 - 28 Mar 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback